how do i do 7r-5r+2=5r-r
Answer by Theo(13342)

equation is:

7r-5r+2=5r-r

combine like terms to get:

2r + 2 = 4r

subtract 2r from both sides of equation to get:

2 = 2r

divide both sides of equation by 2 to get r = 1

substitute in original equation to see if this is good.

original equation is:

7r-5r+2=5r-r

substitute 1 for r to get:

7 - 5 + 2 = 5 - 1 which becomes:

4 = 4 which is true, confirming that r = 1 is good.